OpenStudy (anonymous):

How to find inverse laplace transform of a given function urgent help needed!!!!!!

OpenStudy (anonymous):

2 / ( s^2 + 3s - 4 )

OpenStudy (nikvist):

\[\frac{2}{s^2+3s-4}=\frac{2}{(s-1)(s+4)}=\frac{2/5}{s-1}+\frac{-2/5}{s+4}\] \[L^{-1}\left(\frac{2}{s^2+3s-4}\right)=\frac{2}{5}e^t-\frac{2}{5}e^{-4t}\]

OpenStudy (anonymous):

Where did 2/5 come from ?

OpenStudy (nikvist):

\[\frac{2}{(s-1)(s+4)}=\frac{A}{s-1}+\frac{B}{s+4}\] then solve it,and you will get A=-B=2/5

OpenStudy (anonymous):

Thanks alot. Any general rule for solving these questions ?